the mean and SD of a random sample of size 145 from a large population are 14.3 and 3.4, respectively. give an estimate of the maximum error for estimating u with 42% confidence
Answer by ewatrrr(24785)

Sample of 145: m = 14.3 and SD = 3.4
(100%-42%)/2 = 29%
invNorm(.29) = .5534
ME = .5534(3.4/(sqrt(145))= .16 (nearest one hundredth)
CI: 14.3-ME < u < 14.3 + ME